OpenStudy (anonymous):

Prove that \[\frac{1}{2}+\cos x+\cos2x+\cdots+\cos nx=\frac{\sin\left(n+\dfrac{x}{2}\right)}{2\sin\dfrac{x}{2}}\] (If anyone discovers that this actually isn't an identity, please let me know...)

OpenStudy (fibonaccichick666):

can we use induction?

OpenStudy (fibonaccichick666):

I'm not certain, but I can try here if no one knows, please if you know interrupt me

OpenStudy (anonymous):

Any way you see fit! I haven't tried it myself, so I can't say if it'll work.

OpenStudy (fibonaccichick666):

So we need to show it is true for every n, then prove for n+1.

OpenStudy (fibonaccichick666):

I amthinking we can use identities

OpenStudy (fibonaccichick666):

so what if we work backwards

OpenStudy (fibonaccichick666):

If we look at sin(A+B)=? then go through, it may work

OpenStudy (fibonaccichick666):

ok maybe check this out, I may be out of my league http://math.stackexchange.com/questions/520719/prove-frac12-cosx-cos2x-dots-cosnx-frac-sinn-frac1

OpenStudy (fibonaccichick666):

I feel like proving those hints may not be simple though

OpenStudy (michele_laino):

for inductionn n: for n=1, w have: \[\frac{ 1 }{ 2 }=\frac{ 1 }{ 2 }\] now I suppose that your expression is true for n-1, then for n: \[\frac{ 1 }{ 2 }+\cos x+...+\cos ((n-1)x)+\cos (nx)=\frac{ \sin(n-1/2)x }{ 2 \sin x }+\cos nx=\] \[=\frac{ \sin(n-1/2)x+2 \cos nx \sin (x/2) }{ 2 \sin (x/2) }\] we get our thesis,if we note that \[\sin \left( n-\frac{ 1 }{ 2 } \right)x=\sin nx \cos (x/2)\cos nx \sin (x/2)\]

OpenStudy (michele_laino):

oops for n=0

OpenStudy (fibonaccichick666):

but the question is how do we get that, how can we modify our original into that

OpenStudy (fibonaccichick666):

(I use strong induction ususally, but they are logically equivalent)

OpenStudy (michele_laino):

oops i have made an error of typo: \[\sin \left( n-\frac{ 1 }{ 2 } \right)x=\sin nx \cos(x/2)-\cos nx \sin(x/2)\]

OpenStudy (michele_laino):

we have to subtitute the identity above into the last step of the application of induction principle

OpenStudy (michele_laino):

after that substitution, we get the subsequent expression at numerator: \[\sin nx \cos(x/2)-\cos nx \sin(x/2)+2 \cos nx \sin(x/2)=\] \[=\sin nx \cos (x/2)+\cos nx \sin (x/2)=\sin \left( n+\frac{ 1 }{ 2 } \right)x\]

ganeshie8 (ganeshie8):

wow! never seen an induction version for this before looks very interesting xD Here is an interesting problem that can be solved easily using this identity http://openstudy.com/study#/updates/53ec6330e4b0f30a87d43077

ganeshie8 (ganeshie8):

Alternative : \[\frac{1}{2}+\sum\limits_{k=1}^n \cos(kx) = \frac{1}{2} + \mathfrak{Real}\sum\limits_{k=1}^n e^{ikx} = \cdots \] Notice it is a geometric series with common ratio of \(e^{ix}\) so we can evaluate using partial sum formula and take the real part

OpenStudy (anonymous):

Yup I was shooting for the complex exponential form and using induction on that. Nice work @Michele_Laino !

OpenStudy (michele_laino):

Thank you! @SithsAndGiggles